FEL Upgrade Project Weekly Brief
September 18-22, 2000








Management

We completed preparations for Jefferson Labís annual Science and Technology Peer Review which will be held next week (Sept. 25-27) by SURA and DOE management.  Sam Krinsky from BNL and Max Cornacchia from SSRL will be reviewing FEL related activities.  Steve Benson and Jim Boyce represented the FEL team and gave presentations at the Opto
Southeast meeting held in Charlotte on Sept. 18-19.
 

WBS 3 (Beam Physics)
The following is summary for the weeks of 9/11 and 9/19:

Longitudinal phase space dynamics for the upgrade were investigated and scenarios for nominal operation were established. In addition, a configuration leading to ultra-short bunches (~10-20 fsec rms) was developed. This work is documented in JLAB-TN-00-020, available on line at

http://www.jlab.org/~douglas/FELupgrade/technote/JLABTN00020.pdf

Momentum acceptance measurements were made on the driver; the latest (9/22) indicated the full momentum aperture was only of order 5%, in contrast to the 6+% indicated by FEL operation. Fascinating nonlinear effects (such as T346 induced vertical dispersion) were observed and corrected by orbit ironing; additional work involved implementation of the multi-(quad and sextupole) family compaction and  dispersion management schemes considered necessary in the upgrade.

Initial work on UV  bypass concepts yielded four candidate endloop designs, 2 based on a stair-case (reverse bending with dispersion modulation) geometry and two based on a single step (dispersion modulation only) geometry. Analysis of each in a complete recirculator layout will start shortly.

Beam dynamics and engineering personnel initiated discussion of the energy recovery dump and developed layout options for both upgraded Demo and large aperture "garbage can" style dumps. Once established, these foot-prints will provide guidance for  the beam optics design of, and instrumentation for, this region.
 

WBS 4 (Injector)

We ran the field emission test system up to 120 kV with a 1 cm gap on the nitrogen implanted/SS electrodes.  Currents were very low, a few nanoamps for 12Mv/m.  Set up and started a vacuum bake on the system but there appears to be a problem with the ion pump being able to handle the gas load at an elevated temperature.  Cooling the system down to investigate.

o Light Box preliminary design issues are being discussed. Final concepts discussions are planned.
 

WBS 5 (SRF)

Had an initial vendor visit to start vacuum vessel and space frame fabrication. No outstanding cost, technical, or schedule issues. Vendor is well prepared for contract.

Waveguide and end can procurements continue to make progress.

Signed off top hat design drawings.

Heat shield design package is almost complete and is expected to be finished next week.
 

WBS 6 (RF)

The Waveguides and HPA cabinet for zone 3 were moved into place.

The blower for the zone 4 Low Level RF racks is being installed today.

Both zones need klystrons, RF Control Modules, and a final checkout. Zone 3 also needs circulators.

The specification for the HVPS for the 100 kW klystrons was given to Purchasing today.
 

WBS 8  (I & C)

The requisition has been signed for the 36 3" shielded viewers. Thanks to the mechanical engineering department and to procurement. The delivery schedule calls for first pieces by April and completion by June 2001, this is 3 months ahead of the critical path for girder assembly.  We are still waiting for budgetaries for the BPMs.

The design for the 32 channel VME based sample and hold board is complete and in EECAD for drawings and board layout. We hope to have a prototype to test during the Oct. run. The optical BPM design is proceeding well this should be ready for a prototype surface mount board layout by next week.

The user patch panels and user LAN connections are being worked on and will be complete by the Oct. run. Also the 32 channel pico-motor multplexer is being finished up to allow control of turning mirrors in each of the user labs.
 

WBS 9 (Transport)

o  Details of the steel yoke parts for the GX prototype are complete.  Drawings are in the checking phase.

o  The 3D model the of the GX dipole, using the optimum bump on the pole edge was constructed and ran to completion.  Processing of results continues.

 o Began TOSCA modeling of GW magnet.

o Differential Pumping stations are being modified to accommodate a three inch aperture.
 

WBS 10 (Wiggler)

o The top assembly drawing of the Optical Klystron was completed. Review and sign-off should happen next week. Details of  the optical path has begun.

o   The purchase recquisition for the fabrication of the dispersion section magnet (GS) has been written.
 

WBS 11 (Optics)

This week, our Upgrade efforts concentrated on the following areas: New optical cavity:  Recalculated parameters for a ~ 32m near-concentric cavity.  Began rewriting requirements document, based on new parameters.  The design work on the mirror test stand is at the point where we're working through check prints for the vacuum vessel.  We operated the FEL to further check-out of the on-line optical diagnostics.  We began using the pyroelectric array with the spectrometer in the optical diagnostics room.  We are also testing out detectors, and have been collecting FEL mode quality and transport
properties (i.e., pointing stability) through the EPICS control system, using equipment mounted at the end of the optical transport line.
 

Operations/Commissioning

We operated the FEL this week two days to make beam transport measurements.  This will help determine the energy aperture of the upgrade.
 
 
 
 
 
 

Privacy and Security Notice
Privacy and Security Notice